# Machine Learning Assisted Parameter Tuning on Wavelet Transform Amorphous Radial Distribution Function

### v1

@@ 5.2 Data scaling on WT-RDF++
To test the generality performance of WT-RDF++ compared with WT-RDF, we present the results in Table 4 and Fig. 5. Specifically, we also compare under different data ratios. We observe that the proposed WT-RDF++ outperforms WT-RDF in terms of MAE, FPE, and SPE. This is further supported by the plotting visualization of the results.
@@ 5.4 Cross-dataset evaluation
To evaluate cross-dataset generalization, we used the optimized parameter of WT-RDF++ from the binary data. The models are then evaluated on unseen data, i.e., ternary A, B, C, D, and E. In Fig. 7, a comparison between WT-RDF++ and WT-RDF reveals that WT-RDF++ consistently delivers improved performance. The optimization improves the adaptability of WT-RDF++ to diverse data distributions while maintaining physical consistency. By balancing the physical prior with machine learning tuning, WT-RDF++ captures subtle variations across datasets, resulting in improved generalization and predictive performance on unseen data.
@@ 5.5 Optimization evaluation
In Fig. 8, the proposed tuning strategy exhibits robust effectiveness on multiple datasets. We also report the parameter changes resulting from the optimization. On the basis of the results, combining physics and machine learning yields improved predictions through mutual synergy. This demonstrates the consistency of predictive performance across all datasets. In addition, it is worth mentioning that, in this study, we present the optimal parameters deduced from training on binary data ($\text{Ge}_{0.25}\text{Se}_{0.75}$), as it represents stable and simpler amorphous materials [36, 37, 38, 39].
Figure 9: Numbers of parameter comparison of WT-RDF++ and ML methods.
@@ 5.6 Performance and parameter variability
As training may involve random elements, such as weight initialization, the results can vary with each run. Our models are trained over three runs. In Table 6, we report the errors and parameters of our optimized model, along with the standard deviation to assess the variability. We observe that the parameter values obtained from machine learning assistance are sensible according to the characteristics of the parameters discussed in Section IV-B-1.
@@ 5.7 Parameter efficiency
We evaluate the parameter efficiency of each method. Fig. 9 shows that our WT-RDF++ only has 5 parameters compared to other ML methods, i.e., RBF and LSTM. Specifically, we merely update the parameters $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$. This result suggests the parameter efficiency of our physics-based model. WT-RDF++ achieves stable learning with significantly fewer samples due to its compact parameterization and strong physical inductive biases. This is particularly advantageous in industrial problems where data acquisition is expensive and physically constrained.
@@ 6.2 Why Physics-based model?
In experimental practice, RDF determination often begins with X-ray diffraction to find the expression $S_{R}(q)$ and transform the data based on FT. However, the use of X-ray diffraction often has limitations. Two examples of these limitations are related to the detectable angle range and the resolution of the X-ray diffraction instrument. The angle range directly affects the wave vector $q$, resulting in minimal data availability for FT input. Furthermore, the poor resolution of X-ray diffraction instruments affects the pre-data for extracting $S_{R}(q)$. This, of course, also has a significant impact on the poor resolution of the RDF produced through FT.
Based on these two facts, a physics-based RDF model is needed. The physics-based RDF model developed in this study, WT-RDF+, enables us to predict RDF with good resolution, despite minimal data availability and the resolution of the X-ray diffraction instrument itself. The reason is that through physics, WT-RDF+ has been able to provide an overview of the typical main interactions between atoms and corrections to the interaction terms in the design of its wavelet function. In addition, WT enables the ability to act as an extraordinary mathematical microscope for function analysis on a small (local) scale, as explained in Section 2. In addition to being physics-based, WT-RDF+ excels because its parameter determination is assisted by the backpropagation algorithm of Machine Learning. The ML assistance presented here is intended to guide the determination of model parameters that are more accurate and enhance the predictive capabilities of the existing WT-RDF, rather than building an ML model or a hybrid model.
@@ 7 Conclusion
This research significantly enhances the accuracy and predictive performance of the WT-RDF model, resulting in the improved WT-RDF+. Its effectiveness is demonstrated by more precise predictions of the first and second RDF peaks and the overall RDF trends in binary $\text{Ge}_{0.25}\text{Se}_{0.75}$ (binary) and ($\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})$ 100-x, $x=(5,10,15,20,25)$) amorphous systems. WT-RDF+ also outperforms machine learning models such as RBF and LSTM, especially with limited data. With only 25% of the training data, it reduces the FPE by 98.74% vs. RBF and 98.85% vs. LSTM, and the SPE by 96.40% and 93.15%, respectively. These results highlight WT-RDF+ as a robust tool for advancing technology and manufacturing with amorphous materials. For example, in thin-film phase-change materials (PCM) like Ge–Se, WT-RDF+ enables accurate predictions of structural transitions and thermal properties, facilitating efficient fabrication and reliable assessment of key material characteristics with minimal data.

### v2

@@ 5.2 Data scaling on WT-RDF++
To test the generality performance of WT-RDF++ compared with WT-RDF, we present the results in Table 4 and Fig. 6. Specifically, we also compare under different data ratios to validate the models’ ability to interpolate data. We observe that the proposed WT-RDF++ outperforms WT-RDF in terms of MAE, FPE, and SPE. This is further supported by the plotting visualization of the results.
@@ 5.4 Cross-dataset evaluation
To evaluate cross-dataset generalization, we used the optimized parameter of WT-RDF++ from the binary data. The models are then evaluated on Ag-doped Ge-Se (ternary A, B, C, D, and E). In Fig. 8, a comparison between WT-RDF++ and WT-RDF reveals that WT-RDF++ consistently delivers improved performance. The optimization improves the adaptability of WT-RDF++ to diverse data distributions while maintaining physical consistency. By balancing the physical prior with machine learning tuning, WT-RDF++ captures subtle variations across datasets, resulting in improved reconstructive performance on Ag-doped Ge-Se datasets compared to WT-RDF.
@@ 5.5 Optimization evaluation
In Fig. 9, the proposed tuning strategy exhibits robust effectiveness on multiple datasets. We also report the parameter changes resulting from the optimization. On the basis of the results, combining physics and machine learning yields improved reconstructions through mutual synergy. This demonstrates the consistency of reconstructive performance across all datasets. In addition, it is worth mentioning that, in this study, we present the optimal parameters deduced from training on binary data ($\text{Ge}_{0.25}\text{Se}_{0.75}$), as it represents stable and simpler amorphous materials [35, 36, 37].
@@ 5.6 Performance and parameter variability
As training may involve random elements, such as weight initialization, the results can vary with each run. Our models are trained over three runs. In Table 6, we report the errors and parameters of our optimized model, along with the standard deviation to assess the variability. We observe that the parameter values obtained from machine learning assistance are sensible according to the characteristics of the parameters discussed in Section 3.2.1.
Figure 10: Numbers of parameter comparison of WT-RDF++ and ML methods.
@@ 5.7 Parameter efficiency
We evaluate the parameter efficiency of each method. Fig. 10 shows that our WT-RDF++ only has 5 parameters compared to other ML methods, i.e., Radial Basis Function (RBF) and Long Short-term Memory (LSTM). Specifically, we merely update the parameters $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$. This result suggests the parameter efficiency of our physics-based model. WT-RDF++ achieves stable learning with significantly fewer samples due to its compact parameterization and strong physical inductive biases. This is particularly advantageous in industrial problems where data acquisition is expensive and physically constrained.
Table 7: Comparison of optimization algorithms.
Method MAE $\downarrow$ FPE $\downarrow$ SPE $\downarrow$
L-BFGS-B (w/o $\mathcal{L}_{SL}$) 0.5092 1.3283 0.5398
L-BFGS-B (w/ $\mathcal{L}_{SL}$) 0.7015 0.2166 0.0243
Adam [31] 0.6598 0.0065 0.0199
5.8 Optimization algorithm
As shown in Table 7, we compare the used optimization algorithm (Adam [31]) in this paper with the classical one, i.e., L-BFGS-B. L-BFGS-B [38] is an algorithm solving large nonlinear optimization problems with bounding variable. In this study, we use the same training settings as our approach to train with the L-BFGS-B including the parameter bounding values of $a$ and $K_{f}$, training iterations, two-rounds training, and the use of selective loss function for a fair comparison. For training, the initial parameters are $a=0.590,b=4.600,K_{f}=0.106,\tilde{C}=0.212,\Lambda=-4.700$. On the basis of the result, Adam achieves better performance compared to L-BFGS-B. We conjecture that the $S_{R}(q)$ data is noisy and non-convex. Therefore, stochastic method like Adam is designed to handle such issue as adaptive learning rates and momentum are incorporated, navigating noisy and non-convex landscapes more effectively. It has also been demonstrated that selective loss plays a vital role in enhancing the FPE and SPE results.
@@ 6.2 Why Physics-based model?
In experimental practice, Radial Distribution Function (RDF) determination often begins with X-ray diffraction to find the expression $S_{R}(q)$ and transform the data based on Fourier Transform (FT). However, the use of X-ray diffraction often has limitations. Two examples of these limitations are related to the detectable angle range and the resolution of the X-ray diffraction instrument. The angle range directly affects the wave vector $q$, resulting in minimal data availability for FT input. Furthermore, the poor resolution of X-ray diffraction instruments affects the pre-data for extracting $S_{R}(q)$. This, of course, also has a significant impact on the poor resolution of the RDF produced through FT.
Based on these two facts, a physics-based RDF model is needed. The physics-based RDF model developed in this study, WT-RDF+, enables us to reconstruct RDF with good resolution, despite minimal data availability and the resolution of the X-ray diffraction instrument itself. The reason is that through physics, WT-RDF+ has been able to provide an overview of the typical main interactions between atoms and corrections to the interaction terms in the design of its wavelet function. In addition, WT enables the ability to act as an extraordinary mathematical microscope for function analysis on a small (local) scale, as explained in Section 2. In addition to being physics-based, WT-RDF+ excels because its parameter determination is assisted by the backpropagation algorithm of Machine Learning (ML). The ML assistance presented here is intended to guide the determination of model parameters that are more accurate and enhance the reconstructive capabilities of the existing WT-RDF, rather than building an ML model or a hybrid model.
@@ 7 Conclusion
This research significantly enhances the accuracy and reconstructive performance of the Wavelet Transform Radial Distribution Function (WT-RDF) model by Machine Learning method via learnable parameter ($a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$) optimization, parameter bounding, and selective loss, resulting in the improved WT-RDF+. Its effectiveness is demonstrated by more precise reconstructions of the first and second RDF peaks and the overall RDF trends in binary $\text{Ge}_{0.25}\text{Se}_{0.75}$ (binary) and $\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})$100-x $(x=5,10,15,20,25)$ amorphous systems by requiring a reference $G(r)$ (AIMD data) for calibration. WT-RDF+ also outperforms Machine Learning models such as Radial Distribution Function (RBF) and Long Short-term Memory (LSTM), especially with limited data. With only 25% of the training data, it reduces the FPE by 98.74% vs. RBF and 98.85% vs. LSTM, and the SPE by 96.40% and 93.15%, respectively. These results highlight WT-RDF+ as a robust tool for advancing technology and manufacturing with Ge-Se and Ag-Ge-Se materials.